OpenStudy (anonymous):

In a 45°- 45°- 90° right triangle, the hypotenuse is 9 cm. How long are the legs? In a 45°- 45°- 90° right triangle, the hypotenuse is 9 cm. How long are the legs? @Mathematics

OpenStudy (saifoo.khan):

Hey there! :D

OpenStudy (anonymous):

hey love!

OpenStudy (eyust707):

ok so we have 

OpenStudy (saifoo.khan):

x^2 + x^2 = 9^2 2x^2 = 81 x = 6.36, x = -6.36 take the +ve value.